OpenStudy (anonymous):

12 is what percent of 96?

OpenStudy (luigi0210):

\[\frac{ x }{ 100 }=\frac{ 12 }{ 96 }\]

OpenStudy (anonymous):

96x=100(96) ?

hero (hero):

@GodBlessYou, wrong set up. You can't use 96 twice.

OpenStudy (luigi0210):

cross multiply

OpenStudy (luigi0210):